# You Only Acquire Sparse-channel (YOAS): A Unified Framework for Dense-channel EEG Generation

### v1

@@ IV-C Optimized Generation Paths
We firstly theoretically formulate the problem of generating dense-channel EEG signals from sparse channels by optimizing a set of cross-channel EEG generation tasks. We further introduce the novel unified framework YOAS for dense-channel EEG generation, leveraging prior information on electrode positioning from the International 10-20 system EEG placement, precise bias EEG data for each channel, and general across-channel EEG synthesis paradigms. Additionally, we validate three conjectures of generation pathways (depicted in Fig. 2), contributing to the advancement of dense-channel EEG signal generation methodologies and pioneering new frontiers in this field.
The proposed YOAS currently falls within the realm of data-driven statistical models, where parameters $L$ and $P$ in Data preparation stage (III-B) of YOAS are empirically set, and their correlation with physical mechanisms remains to be investigated in future research. Also, the threshold values for outlier detection are predominantly determined empirically, thus, future work may explore the adaptive detection of outliers in Data Preprocessing stage (III-C) of YOAS, drawing inspiration from metrics such as the Interquartile Range (IQR) [35].
Our innovative approach revolutionizes the construction of dense-channel EEG systems, promising broader applications. By optimizing EEG systems to support more channels with fewer resources, we potentially significantly reduce the cost of high-density electrode arrays while maintaining its performance. This breakthrough enhances EEG accessibility in healthcare, consumer electronics, and neuroscience.

### v2

@@ V Discussion and Conclusion
We first theoretically formulate the problem of generating dense-channel EEG signals from sparse channels by optimizing a set of cross-channel EEG generation tasks. Then, We introduce the novel unified framework YOAS for dense-channel EEG generation, leveraging prior information on electrode positioning from the International 10-20 system EEG placement, precise bias EEG data for each channel, and general across-channel EEG synthesis paradigms. Additionally, we validate three conjectures of generation pathways (depicted in Fig. 2), contributing to the advancement of dense-channel EEG signal generation methodologies and pioneering frontiers in the field.
The proposed YOAS currently falls within the realm of data-driven statistical models, where threshold values for outlier detection are predominantly determined empirically. Future work may explore the adaptive detection of outliers in the Data Preparation stage of YOAS, drawing inspiration from metrics such as the Interquartile Range (IQR) [66]. Furthermore, parameters $L$ and $P$ are empirically set, and their correlation with physical mechanisms remains to be investigated in future research.
    Our innovative approach revolutionizes the construction of dense-channel EEG systems by developing advanced artificial intelligence models, promising broader applications. By optimizing EEG systems to support more channels with fewer resources, we potentially significantly reduce the cost of high-density electrode arrays while maintaining performance. This breakthrough enhances EEG accessibility in healthcare, consumer electronics, and neuroscience.